# Sufficient conditions for strong discrete maximum principles in finite element solutions of linear and semilinear elliptic equations

### v1

@@ 1. Introduction
The main contribution in this article is to provide a novel technique for proving a global strong DMP (sDMP) for several classes of elliptic equations. The main ingredient consist of extending the sDMPs from macroelements to the entire domain using a connectivity argument; nonpositivity of the off-diagonal entries of the stiffness matrix is not assumed to hold everywhere. The general theorems are applied to linear elliptic equations that include “defects” (edges that lead to positive off-diagonal entries in the stiffness matrix), degenerate meshes, as well as semi-linear elliptic equations.
@@ 2.1. The continuous model problems and their maximum principles
for some constants $0<\underline{d}\leq\overline{d}$. Assume the reaction term $c:D\times\mathbb{R}\to\mathbb{R}$ satisfies the following conditions:
(2.3) $\forall x\in D,\ c(x,0)=0\ \ \mathrm{and}\ \ c(x,\cdot)\ \ \mathrm{is\ nondecreasing;}$
@@
for some nonnegative function $c\in L^{\infty}(D)$. Additional smoothness of the coefficients $a_{ij}$ and the function $c$ is usually imposed. For a function $u$ we define $u^{+}=\max(u,0)$, and $u^{-}=-\min(u,0)$. Note that $u^{+},u^{-}\geq 0$ and $u=u^{+}-u^{-}$.
The continuous problem (2.1a)-(2.1b) is known [10, 7] to satisfy the following strong maximum principles:
@@ 2.2. Finite element discretization
A set $E\subseteq D$ is called a discrete subdomain (of $D$) if $E$ is a union of triangles of ${\mathcal{T}}_{h}$. When solving the problem on a discrete subdomain $E$, then $E$ must replace $D$ everywhere in (2.12). Due to the potential nonlinearity of the reaction term $c$ in the second argument, the term $\left(c(\cdot,u_{h}^{0}+u_{h}^{b}),v\right)_{D}$ is replaced with a cubature, preferably one that involves only function values at the vertices. This is equivalent to interpolating the aforementioned term prior to integration. More precisely, (2.12) can be solved in practice using the following formulation:
@@
The formulations (2.12) and (2.13) are certainly equivalent if $c(x,u)=\tilde{c}u$ with $\tilde{c}\geq 0$ constant, but in general thet are not. In preparation for the results in Section 3, we highlight the following localization property of the finite element solutions:
@@ 3.1. Abstract solution operators
Denote $V^{h}(\partial D)=\{\varphi|_{\partial D}\>:\>\varphi\in V^{h}(\overline{D})\}$. We say that $f\in(V_{0}^{h}(\overline{D}))^{*}$ is nonnegative (or nonpositive, respectively), if $\left<f,\varphi_{i}\right>_{D}\geq 0$ (respectively, $\left<f,\varphi_{i}\right>_{D}\leq 0$), for all nodal basis functions $\varphi_{i}$, $i=1,\dots,N$. Furthermore, if $E\subseteq D$ is a discrete subdomain, we denote by $f_{E}$ the natural restriction of $f$ to $(V_{0}^{h}(\overline{E}))^{*}$, which is defined by $\left<\varphi,f_{E}\right>=\left<\varphi^{D},f\right$>, where $\varphi^{D}$ is the extension with 0 of $\varphi\in V_{0}^{h}(\overline{E})$ to $\overline{D}$. Cleary, if $f$ is nonnegative/nonpositive, then $f_{E}$ is also nonnegative/nonpositive, respectively. The objects for which we prove DMPs are the following abstract solution operators.
@@ 3.2. Main results
Let $({\mathcal{S}}^{E}_{h})_{E\subseteq D}$ be a consistent family of solution operators, and the triangulation ${\mathcal{T}}_{h}$ be so that the graph of the interior vertices of is connected, and that every boundary vertex is adjacent to an interior vertex.
@@
Let $({\mathcal{S}}^{E}_{h})_{E\subseteq D}$ be a consistent family of solution operators, and the triangulation ${\mathcal{T}}_{h}$ be so that the graph of the interior vertices of is connected, and that every boundary vertex is adjacent to an interior vertex. Let $X$ stand for any of the symbols $A$, $B$, or $R$. Assume that there exists a finite set of discrete subdomains $(E_{i})_{i\in I}$ so that
@@ 3.3. Matrix form of sDMP-A and sDMP-B for linear elliptic equations
To prove the reverse implication, assume ${\mathcal{S}}^{D}_{h}$ satisfies sDMP-A. For an internal node $i\in\alpha$ define the source $f^{(i)}$ so that $\left<f,\varphi_{j}\right>=\delta_{ij}$, i.e., $f^{(i)}$ is represented by the basis element ${\mathbf{e}}_{i}$; let $u_{h}^{b}\equiv 0$. Due to sDMP-A, the vector ${\mathbf{u}}^{(i)}$ representing the solution $u_{h}^{(i)}$ satisfies
@@
$({\mathbf{A}}_{\alpha\alpha}+{\mathbf{A}}_{\alpha\alpha})^{-1}{\mathbf{F}}_{\alpha}\geq 0.$
@@
Let $({\mathcal{S}}^{E}_{h})_{E\subseteq D}$ be a consistent family of solution operators associated with the discrete linear variational inequality (2.12). We assume the triangulation ${\mathcal{T}}_{h}$ is so that the graph of the interior vertices is connected, and that every boundary vertex is adjacent to an interior vertex. Furthermore, assume that there exists a finite set of discrete subdomains $(E_{i})_{i\in I}$ so that $Int(D)=\cup_{i\in I}Int(E_{i})$, and denote by $\alpha_{i}$ the set of interior vertices of $E_{i}$ and by $\beta_{i}$ the set of boundary vertices of $E_{i}$.
@@ 3.4. Sufficient conditions for a weak discrete maximum principle
Now we construct the family of consistent solution operators, namely we consider the following sequence of discrete variational inequalities: given a discrete subdomain $E\subseteq D$, find $u_{h}\in V^{h}(\overline{E})$ of the form $u_{h}=u_{h}^{0}+u_{h}^{b}$ with $u_{h}^{0}\in V_{0}^{h}(\overline{E})$, so that
@@ 4.1. Connection to the classical results
Let $E\subseteq D$ a connected discrete subdomain. For every vertex $P_{i}$ interior to $E$ we consider the set $E_{i}=\bigcup\{T\in{\mathcal{T}}_{h}\ :\ P_{i}\in T\}$, which we call the star around $P_{i}$. The boundary vertices of $E_{i}$ are the adjacent vertices of $P_{i}$, which is the only interior node of $E_{i}$. In the terminology of Lemma 3.6, $\alpha=\{i\}$ and $\beta=\{j\ :\ {\mathbf{A}}_{ij}<0$, and the conditions (3.8) are satisfied in a trivial manner. Hence, ${\mathcal{S}}^{E_{i}}_{h}$ satisfies sDMP-A. Corollary 3.5 implies that ${\mathcal{S}}^{E}_{h}$ satisfies sDMP-A. ∎
@@ 4.2. The angle condition
(4.1) ${\mathbf{A}}_{ij}=a_{D}(\varphi_{j},\varphi_{i})=\int_{{\mathcal{T}}_{1}\cup{\mathcal{T}}_{2}}\nabla\varphi_{i}\cdot\nabla\varphi_{i}=-\frac{\sin(\theta_{1}+\theta_{2})}{2\sin\theta_{1}\sin\theta_{2}}.$
@@
$Q(f)=\frac{\sigma({\mathcal{T}})}{3}\sum_{\ell\in\{i,j,k\}}f(M_{\ell})$
is exact for quadratics, where $M_{i},M_{j},M_{k}$ are the edge midpoints (see Fig. 1, right). The contributions of $\mathcal{T}$} to the $i^{\mathrm{th}}$ row of the mass matrix are
(4.4) $\int_{{\mathcal{T}}}\varphi_{i}^{2}$ $=$ $\frac{\sigma({\mathcal{T}})}{3}\sum_{\ell\in\{i,j,k\}}\varphi^{2}_{i}(M_{\ell})=\frac{\sigma({\mathcal{T}})}{6}=\frac{h_{i}^{2}}{12(\cot\theta_{j}+\cot\theta_{k})},$
(4.5) $\int_{{\mathcal{T}}}\varphi_{i}\varphi_{j}$ $=$ $\frac{\sigma({\mathcal{T}})}{3}\sum_{\ell\in\{i,j,k\}}\varphi_{i}(M_{\ell})\varphi_{j}(M_{\ell})=\frac{\sigma({\mathcal{T}})}{12}=\frac{h_{k}^{2}}{24(\cot\theta_{i}+\cot\theta_{j})},$
@@
$\sigma({\mathcal{T}})=\frac{h_{i}^{2}}{2(\cot\theta_{j}+\cot\theta_{k})}.$
@@ 4.3. Defects related to boundary mesh properties
Case 1: $\theta>\pi/2$ (Fig. 2 – right). We first note that the boundary values imposed at the vertices $P_{n+1}=(1,0)$ and $P_{k}=(\cos\theta,\sin\theta)$, with $k=1+n(n+1)$, do not influence the solution in the interior; that is, $u^{0}_{h}$ is the same regardless of the values $u^{b}_{h}(P_{n+1})$ and $u^{b}_{h}(P_{k})$. This shows $S^{D}_{h}$ does not satisfy sDMP-A, since attaining a minimum for $u_{h}^{0}$ in the interior does not imply $u_{h}=u^{0}_{h}+u^{b}_{h}$ is constant. However, if $\tilde{D}$} denotes the domain $D$ from which we remove the triangles containing the vertices $P_{n+1}$ and $P_{k}$ (shaded in Fig. 2 – right), then $S^{\tilde{D}}_{h}$ satisfies sDMP-A, since all the angles are acute, and every boundary node is connected to an internal node. Furthemore, if we now add back the triangles to the domain, then for any nonnegative $f$ and discrete boundary function $u^{b}_{h}\in V^{h}(\partial D)$, the interior part of the solution $u^{0}_{h}$ satisfies
@@ 5.1. A class of meshes with defects
Figure 5. Smallest values of ${\mathbf{A}}^{-1}_{1},\dots,{\mathbf{A}}^{-1}_{4}$ as functions of $\theta$; note that they become positive as $\theta\to\pi/2$.
@@ 5.2. More complicated meshes
Let $k\in\mathbb{N}$ and $\theta$ so that $\max_{\ell=1}^{k}\theta_{\ell}<\theta<\frac{\pi}{2}$. Consider a uniform partition of $D=D(\theta)$ in $n\times n$ rhombuses, with each small rhombus being divided along either its long diagonal or its short diagonal, and let $h$ be the mesh size. We assume that each interior vertex $P$ lies in the interior of a mesh that is a scaled version of $G_{\ell}(\theta)$ for some $1\leq\ell\leq k$, or the star around $P$ has only acute angles. (Examples of such meshes are given in Figures 6-7). Let $({\mathcal{S}}^{E}_{h})_{E\subseteq D}$ be the consistent family of solution operators associated with the discrete linear variational inequality (2.12). Then the following hold:
@@
The argument for (ii) lies in the scaling of the mass matrix compared to that of the stiffness matrix. Assume a point $P$ lies in the interior of a patch $F_{P}$ that is a rescaled version of $G_{\ell}(\theta)$, as above. Denote by ${\mathbf{M}}_{\ell}={\mathbf{M}}_{\ell}(\theta)$ the mass matrix associated with the interior vertices of the reference mesh $G_{\ell}(\theta)$. Due to scaling (see (4.5)), the associated mass matrix for $F_{P}$ is $Kh^{2}{\mathbf{M}}_{\ell}$, with $K$ a dimensionless constant depending on $\theta$ and $\ell$. Similarly, if ${\mathbf{M}}^{b}_{\ell}={\mathbf{A}}^{b}_{\ell}(\theta)$ is the mass matrix connecting interior and boundary entries on $G_{\ell}(\theta)$, then the analogous matrix on $F_{P}$ is $Kh^{2}{\mathbf{M}}^{b}_{\ell}$. Hence, the interior and boundary reaction matrices ${\mathbf{C}}_{P}$ and ${\mathbf{C}}^{b}_{P}$, respectively, associated with $F_{P}$ (see (3.7)) satisfy
@@
If, instead, the star $E_{P}$ around $P$ has only acute angles, similar inequalities hold if we replace $F_{P}$ with $E_{P}$. By Corollary 3.9(ii), satisfies ${\mathcal{S}}^{\tilde{D}}_{h}$ satisfies sDMP-B. The argument shown in Example 4.4 can be replicated to show that when adding back the two triangles at the horizonal extremes of $\tilde{D}$, the global solution operator ${\mathcal{S}}^{{D}}_{h}$ satisfies wDMP-B. ∎
@@
Figure 9. Uniform three-line mesh with an embedded, irregularly divided triangle one layer away from the boundary. The finite element solver for the Poisson equation satisfies the wDMP-A on this mesh, but fail to do so if the triangle is placed at the boundary.
@@
There exists a singular matrix ${{\mathbf{T}}_{0}}>{\mathbf{0}}$ so that
@@
The proof of Theorem 5.4 is technical, but elementary; it is given in Appendix A, where also the precise value for ${\mathbf{T}}_{0}$ is also shown. The theoretical result above was validated numerically.

### v2

@@ 1. Introduction
The main contribution in this article is to provide a novel technique for proving a global strong DMP (sDMP) for several classes of elliptic equations. The main ingredient consists of extending the sDMPs from macroelements to the entire domain using a connectivity argument; nonpositivity of the off-diagonal entries of the stiffness matrix is not assumed to hold everywhere. The general theorems are applied to linear elliptic equations that include “defects” (edges that lead to positive off-diagonal entries in the stiffness matrix), degenerate meshes, as well as semi-linear elliptic equations.
@@ 2.1. The continuous model problems and their maximum principles
for some constants $0<\underline{a}\leq\overline{a}$. Assume the reaction term $c:D\times\mathbb{R}\to\mathbb{R}$ satisfies the following conditions:
(2.3) $\forall x\in D,\ c(x,0)=0,\ \ \mathrm{and}\ \ c(x,\cdot)\ \ \mathrm{is\ nondecreasing;}$
(2.4) ${\color[rgb]{0,0,0}c(\cdot,0)\in L^{\frac{p}{2}}(D)\ \mathrm{for\ some\ }p>d;}$
@@
for some nonnegative function $c\in L^{\infty}(D)$. For existence, uniqueness, and regularity results for (2.1a)-(2.1b) see [28] and the references therein.
For a function $u$ we define $u^{+}=\max(u,0)$, and $u^{-}=-\min(u,0)$. Note that $u^{+},u^{-}\geq 0$ and $u=u^{+}-u^{-}$. The continuous problem (2.1a)-(2.1b) is known [12, 9] to satisfy the following strong maximum principles:
@@ 2.2. Finite element discretization
A set $E\subseteq D$ is called a discrete subdomain (of $D$) if $E$ is a union of simplices of ${\mathcal{T}}_{h}$. When solving the problem on a discrete subdomain $E$, then $E$ must replace $D$ everywhere in (2.13). Due to the potential nonlinearity of the reaction term $c$ in the second argument, the term $\left(c(\cdot,u_{h}^{0}+u_{h}^{b}),v\right)_{D}$ is replaced with a cubature, preferably one that involves only function values at the vertices. This is equivalent to interpolating the aforementioned term prior to integration. More precisely, (2.13) can be solved in practice using the following formulation:
@@
The formulations (2.13) and (2.14) are certainly equivalent if $c(x,u)=\tilde{c}u$ with $\tilde{c}\geq 0$ constant, but in general they are not. Cf. [5], the discrete variational problem (2.11) is well-posed. In preparation for the results in Section 3, we highlight the following localization property of the finite element solutions:
@@ 3.1. Abstract solution operators
Denote $V^{h}(\partial D)=\{\varphi|_{\partial D}\>:\>\varphi\in V^{h}(\overline{D})\}$. We say that $f\in(V_{0}^{h}(\overline{D}))^{*}$ is nonnegative (or nonpositive, respectively), if $\left<f,\varphi_{i}\right>_{D}\geq 0$ (respectively, $\left<f,\varphi_{i}\right>_{D}\leq 0$), for all nodal basis functions $\varphi_{i}$, $i=1,\dots,N$. Furthermore, if $E\subseteq D$ is a discrete subdomain, we denote by $f_{E}$ the natural restriction of $f$ to $(V_{0}^{h}(\overline{E}))^{*}$, which is defined by $\left<\varphi,f_{E}\right>=\left<\varphi^{D},f\right$>, where $\varphi^{D}$ is the extension with 0 of $\varphi\in V_{0}^{h}(\overline{E})$ to $\overline{D}$. Clearly, if $f$ is nonnegative/nonpositive, then $f_{E}$ is also nonnegative/nonpositive, respectively. The objects for which we prove DMPs are the following abstract solution operators.
@@ 3.2. Main results
Let $({\mathcal{S}}^{E}_{h})_{E\subseteq D}$ be a consistent family of solution operators, and the triangulation ${\mathcal{T}}_{h}$ be so that the graph of the interior vertices is connected, and that every boundary vertex is adjacent to an interior vertex.
@@
Let $({\mathcal{S}}^{E}_{h})_{E\subseteq D}$ be a consistent family of solution operators, and the triangulation ${\mathcal{T}}_{h}$ be so that the graph of the interior vertices is connected, and that every boundary vertex is adjacent to an interior vertex. Let $X$ stand for any of the symbols $A$, $B$, or $R$. Assume that there exists a finite set of discrete subdomains $(E_{i})_{i\in I}$ so that
@@ 3.3. Matrix form of sDMP-A and sDMP-B for linear elliptic equations
To prove the reverse implication, assume ${\mathcal{S}}^{D}_{h}$ satisfies sDMP-A. For an internal node $i\in\alpha$ define the source $f^{(i)}$ so that $\left<f^{(i)},\varphi_{j}\right>=\delta_{ij}$, i.e., $f^{(i)}$ is represented by the basis element ${\mathbf{e}}_{i}$; let $u_{h}^{b}\equiv 0$. Due to sDMP-A, the vector ${\mathbf{u}}^{(i)}$ representing the solution $u_{h}^{(i)}={\mathcal{S}}^{D}_{h}(f^{(i)},0)$ satisfies
@@
We remark that the vector ${\mathbf{u}}^{(i)}$ in the proof of the reverse implication is the vector representation of discrete Green’s function with Dirac impulse forcing $f(x)=\delta(x-P_{i})$. Hence, the matrix ${\mathbf{A}}_{\alpha\alpha}^{-1}$ represents the discrete Green’s function, and the first of the conditions in (3.8) states that the discrete Green’s function be positive. Oftentimes this is regarded as being equivalent to the DMP. However, we note that this form of the DMP only applies to zero–Dirichlet boundary conditions.
@@
$({\mathbf{A}}_{\alpha\alpha}+{\mathbf{C}}_{\alpha\alpha})^{-1}{\mathbf{F}}_{\alpha}\geq 0.$
@@
Let $({\mathcal{S}}^{E}_{h})_{E\subseteq D}$ be a consistent family of solution operators associated with the discrete linear variational problem (2.13). We assume the triangulation ${\mathcal{T}}_{h}$ is so that the graph of the interior vertices is connected, and that every boundary vertex is adjacent to an interior vertex. Furthermore, assume that there exists a finite set of discrete subdomains $(E_{i})_{i\in I}$ so that $Int(D)=\cup_{i\in I}Int(E_{i})$, and denote by $\alpha_{i}$ the set of interior vertices of $E_{i}$ and by $\beta_{i}$ the set of boundary vertices of $E_{i}$.
@@ 3.4. Sufficient conditions for a weak discrete maximum principle
Now we construct the family of consistent solution operators, namely we consider the following sequence of discrete variational problems: given a discrete subdomain $E\subseteq D$, find $u_{h}\in V^{h}(\overline{E})$ of the form $u_{h}=u_{h}^{0}+u_{h}^{b}$ with $u_{h}^{0}\in V_{0}^{h}(\overline{E})$, so that
@@ 4.1. Connection to the classical results
Let $E\subseteq D$ a connected discrete subdomain. For every vertex $P_{i}$ interior to $E$ we consider the set $E_{i}=\bigcup\{T\in{\mathcal{T}}_{h}\ :\ P_{i}\in T\}$, which we call the star around $P_{i}$. The boundary vertices of $E_{i}$ are the adjacent vertices of $P_{i}$, which is the only interior node of $E_{i}$. In the terminology of Lemma 3.6, $\alpha=\{i\}$ and $\beta=\{j\ :\ {\mathbf{A}}_{ij}<0\}$, and the conditions (3.8) are satisfied in a trivial manner. Hence, ${\mathcal{S}}^{E_{i}}_{h}$ satisfies sDMP-A. Corollary 3.5 implies that ${\mathcal{S}}^{E}_{h}$ satisfies sDMP-A. ∎
@@ 4.2. The angle condition
(4.1) ${\mathbf{A}}_{ij}=a_{D}(\varphi_{j},\varphi_{i})={\color[rgb]{0,0,0}\int_{{\mathcal{T}}_{1}\cup{\mathcal{T}}_{2}}\nabla\varphi_{i}\cdot\nabla\varphi_{j}}=-\frac{\sin(\theta_{1}+\theta_{2})}{2\sin\theta_{1}\sin\theta_{2}}.$
@@
$Q(f)=\frac{\mu({\mathcal{T}})}{3}\sum_{\ell\in\{i,j,k\}}f(M_{\ell})$
is exact for quadratics, where $M_{i},M_{j},M_{k}$ are the edge midpoints (see Fig. 1, right), and $\mu({\mathcal{T}})$ denotes the area of $\mathcal{T}$}. The contributions of $\mathcal{T}$} to the $i^{\mathrm{th}}$ row of the mass matrix are
(4.4) $\int_{{\mathcal{T}}}\varphi_{i}^{2}$ $=$ $\frac{\mu({\mathcal{T}})}{3}\sum_{\ell\in\{i,j,k\}}\varphi^{2}_{i}(M_{\ell})=\frac{\mu({\mathcal{T}})}{6}=\frac{h_{i}^{2}}{12(\cot\theta_{j}+\cot\theta_{k})},$
(4.5) $\int_{{\mathcal{T}}}\varphi_{i}\varphi_{j}$ $=$ $\frac{\mu({\mathcal{T}})}{3}\sum_{\ell\in\{i,j,k\}}\varphi_{i}(M_{\ell})\varphi_{j}(M_{\ell})=\frac{\mu({\mathcal{T}})}{12}=\frac{h_{k}^{2}}{24(\cot\theta_{i}+\cot\theta_{j})},$
@@
$\mu({\mathcal{T}})=\frac{h_{i}^{2}}{2(\cot\theta_{j}+\cot\theta_{k})}.$
@@ 4.3. Defects related to boundary mesh properties
Case 1: $\theta>\pi/2$ (Fig. 2 – right). We first note that the boundary values imposed at the vertices $P_{n+1}=(1,0)$ and $P_{k}=(\cos\theta,\sin\theta)$, with $k=1+n(n+1)$, do not influence the solution in the interior; that is, $u^{0}_{h}$ is the same regardless of the values $u^{b}_{h}(P_{n+1})$ and $u^{b}_{h}(P_{k})$. This shows $S^{D}_{h}$ does not satisfy sDMP-A, since attaining a minimum for $u_{h}^{0}$ in the interior does not imply $u_{h}=u^{0}_{h}+u^{b}_{h}$ is constant. However, if $\tilde{D}$} denotes the domain $D$ from which we remove the triangles containing the vertices $P_{n+1}$ and $P_{k}$ (shaded in Fig. 2 – right), then $S^{\tilde{D}}_{h}$ satisfies sDMP-A, since all the angles are acute, and every boundary node is connected to an internal node. Furthermore, if we now add back the triangles to the domain, then for any nonnegative $f$ and discrete boundary function $u^{b}_{h}\in V^{h}(\partial D)$, the interior part of the solution $u^{0}_{h}$ satisfies
@@ 5.1. A class of meshes with defects
Figure 5. The smallest values of ${\mathbf{A}}^{-1}_{1}(\theta),\dots,{\mathbf{A}}^{-1}_{4}(\theta)$ are shown as functions of $\theta$ (on the $x$-axis). Note that each of the four functions has a root $\theta_{k}<\pi/2$ so that it is positive for $\theta_{k}<\theta<\pi/2$.
@@ 5.2. More complicated meshes
Let $k\in\mathbb{N}$ and $\theta$ so that $\max_{\ell=1}^{k}\theta_{\ell}<\theta<\frac{\pi}{2}$. Consider a uniform partition of $D=D(\theta)$ in $n\times n$ rhombuses, with each small rhombus being divided along either its long diagonal or its short diagonal, and let $h$ be the mesh size. We assume that each interior vertex $P$ lies in the interior of a mesh that is a scaled version of $G_{\ell}(\theta)$ for some $1\leq\ell\leq k$, or the star around $P$ has only acute angles. (Examples of such meshes are given in Figures 6-7). Let $({\mathcal{S}}^{E}_{h})_{E\subseteq D}$ be the consistent family of solution operators associated with the discrete linear variational problem (2.13). Then the following hold:
@@
The argument for (ii) lies in the scaling of the mass matrix compared to that of the stiffness matrix. Assume a point $P$ lies in the interior of a patch $F_{P}$ that is a rescaled version of $G_{\ell}(\theta)$, as above. Denote by ${\mathbf{M}}_{\ell}={\mathbf{M}}_{\ell}(\theta)$ the mass matrix associated with the interior vertices of the reference mesh $G_{\ell}(\theta)$. Due to scaling (see (4.5)), the associated mass matrix for $F_{P}$ is $Kh^{2}{\mathbf{M}}_{\ell}$, with $K$ a dimensionless constant depending on $\theta$ and $\ell$. Similarly, if ${\mathbf{M}}^{b}_{\ell}={\mathbf{M}}^{b}_{\ell}(\theta)$ is the mass matrix connecting interior and boundary entries on $G_{\ell}(\theta)$, then the analogous matrix on $F_{P}$ is $Kh^{2}{\mathbf{M}}^{b}_{\ell}$. Hence, the interior and boundary reaction matrices ${\mathbf{C}}_{P}$ and ${\mathbf{C}}^{b}_{P}$, respectively, associated with $F_{P}$ (see (3.7)) satisfy
@@
If, instead, the star $E_{P}$ around $P$ has only acute angles, similar inequalities hold if we replace $F_{P}$ with $E_{P}$. By Corollary 3.9(ii), satisfies ${\mathcal{S}}^{\tilde{D}}_{h}$ satisfies sDMP-B. The argument shown in Example 4.4 can be replicated to show that when adding back the two triangles at the horizontal extremes of $\tilde{D}$, the global solution operator ${\mathcal{S}}^{{D}}_{h}$ satisfies wDMP-B. ∎
@@
Figure 9. Uniform three-line mesh with an embedded, irregularly divided triangle one layer away from the boundary. The finite element solver for the Poisson equation satisfies the wDMP-A on this mesh, but fails to do so if the triangle is placed at the boundary.
@@
There exists a singular, rank-3 deficient matrix ${{\mathbf{T}}_{0}}>{\mathbf{0}}$ so that
@@
The proof of Theorem 5.4 is technical, but elementary; it is given in Appendix A, where also the precise value for ${\mathbf{T}}_{0}$ is also shown. The theoretical result above was validated numerically. With regards to DMPs, the most remarkable part of Theorem 5.5 is the positivity of ${\mathbf{T}}_{0}$ as $\alpha\to 0$. The existence of the limit as a bounded and singular matrix is related to the behavior of discrete Green’s functions on degenerate meshes. In order to better illustrate this behavior we provide a simpler case in Appendix B, which emerged from an example discussed in [8] where the DMP holds on certain meshes violating the angle condition.